A farmer wishes to raise 300 chick but estimates that only 60% of the eggs that he puts into the incubator will hatch. How many eggs should be put into the incubator?
Answer by richwmiller(17219)

.6*x=300
6/10*x=300
6x=3000
x=500
check
.6=3/5
3/5*500=300
ok